When 11.8 of C8H18 underwent a combustion reaction with 8.84g of O2 it produced 1.15g of H2O. Balance the reaction and answer the following questions. What is the limiting reactant?


1
Expert's answer
2021-11-19T07:35:04-0500

2C8H18 + 25O2 = 18H2O + 16CO2

n = m / M

M (H2O) = 18 g/mol

M (C8H18) = 114.2 g/mol

M (O2) = 32 g/mol

n (H2O) = 1.15 / 18 = 0.064 mol

According to the reaction, in order to form this amount of water the required amount of incoming reactants is:

n (C8H18)required = 2/18 x n (H2O) = 2/18 x 0.064 = 0.007 mol

n (O2)required = 25/18 x n (H2O) = 25/18 x 0.064 = 0.089 mol


The actual amount of incoming reactants is:

n (C8H18)actual = 11.8/114.2 = 0.1 mol

n (O2)actual = 8.84 / 32 = 0.28 mol


Due to this, both incoming reactants are in excess.